OpenStudy (anonymous):

x2+6x-8=0 soleve by completeing the square ! please help im doing it on the exam and i dont know this one !

OpenStudy (anonymous):

thnks for helping!

OpenStudy (campbell_st):

you need to find the value of (6/2)^2 rewrite the equation \[x^2 + 6x = 8\] then add (6/2)^2 to both sides \[x^2 + 6x + 9 = 8 + 9\] which can be factorised \[(x +3)^2 = 17\] now solve for x \[x + 3 = \pm \sqrt{17}\] then \[x = - 3 \pm \sqrt{17}\]

OpenStudy (anonymous):

so x=-3+-sqr root 17 ?

OpenStudy (anonymous):

thats the answer?

OpenStudy (campbell_st):

yes... hope it helps

OpenStudy (anonymous):

thank youu!!!!